OpenStudy (anonymous):

for a certain quadratic equation ax^2+bx+c=0, the 2 solutions are x= 3\4 and x=-2/5. What are the factors of a^2+bx+c?

OpenStudy (zzr0ck3r):

note for the first one, if we get everything equal to 0 we get x=-2/5 implies 5x=-2 imples 5x+2=0 so one of the factors is (5x+2) can you find the other?

OpenStudy (anonymous):

would it be 4x-3 and 5x +2 ? @skullpatrol & @zzr0ck3r

OpenStudy (zzr0ck3r):

correct

OpenStudy (anonymous):

thank you I understanad really well what you did ! @zzr0ck3r

OpenStudy (zzr0ck3r):

note that the whole point of doing this is so that we can get ab=0 then we are assure that either a = 0 or b = 0 or both =0. 0 is the ONLY number that has this property so we are trying to get things in the form ax^2+bx+c in the form of of two things multiplied together, then we can play with the smaller linear problems which we are guaranteed to get solutions.